# Title of the Paper

### v2

@@ 4.2 Aggregate Welfare
Finally, when the interaction network is regular, the welfare effect of public information increases with agents’ common out-degree. [^14] Suppose $d_{i}^{out}=d$ for all $i\in N$. Then, the term $c_{i}^{2}(\gamma,G)-\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)$ simplifies to $(1+d-2\gamma d^{2})/(1-\gamma d)^{2}$. Differentiating this ratio with respect to $d$ yields $((1-\gamma d)+2\gamma(1-d))/(1-\gamma d)^{3}$, which is strictly positive for all $d\in[0,1)$. Another implication of Proposition 3 (and Example 2) is that this monotonicity does not necessarily hold for more general interaction networks. The following proposition provides a class of information structures under which the welfare effect of public information can decrease as agents’ coordination motives become stronger.
Proposition 4.
[…]
4.3 Implications on Endogenous Information Sharing
@@
The two-stage game is solved by backward induction. The analysis requires comparing equilibrium payoffs under public sharing with those under private retention of the signal. As derived in Section 6, the equilibrium actions in the subgame following a choice to keep the signal private are
@@
Let $\mathcal{J}(\gamma,l,m)\equiv\{(\alpha,\beta)\in[0,1)^{2}:f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}<mf(\alpha,\beta,\gamma)+l(1-\beta(2\gamma\beta-1))\}$ be the set of interaction parameters $(\alpha,\beta)$ that admit inefficiently low equilibrium information sharing, given $\gamma$, $l$, and $m$. Because agent 1’s preference over information structures is “more aligned” with peripheral agents than with core agents, $\mathcal{J}(\gamma,l,m)$ shrinks as the number of peripheral agents relatively increases, i.e., $\mathcal{J}(\gamma,n-m^{\prime},m^{\prime})\subseteq\mathcal{J}(\gamma,n-m,m)$ for $m\leq m^{\prime}$ (see Figure 4). [^17] The formal argument is as follows. Fix $\gamma$ and take any $(\alpha,\beta)$ that satisfies $f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}$. The right-hand side of (25) is linear in $m$ with slope $f(\alpha,\beta,\gamma)-(1-\beta(2\gamma\beta-1))$ by substituting $l=n-m$. This slope is bounded from above by $f(\alpha,\beta,\gamma)-1+2\gamma\beta^{2}-\beta$ $\leq f(\alpha,\beta,\gamma)-1+2\gamma\beta-\beta$ $\leq f(\alpha,\beta,\gamma)-1+2\gamma\beta-\gamma\beta$ $=f(\alpha,\beta,\gamma)-1+\gamma\beta$ $<0.$ The last strict inequality follows from the initial assumption. Thus, the right-hand side of (25) strictly decreases with $m$ whenever $f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}$, which proves the claim.
@@ 5 Conclusion
This paper studies how payoff heterogeneity affects the social value of information in beauty contest games where a network of coordination motives represents agents’ interactions. The analysis provides a geometric characterization of interaction networks under which public information provision can harm aggregate welfare. A key mechanism identified is that agents with relatively low Katz-Bonacich centrality may be harmed by public information because their highly central neighbors, by excessively responding to the public signal, distort their own actions relative to the payoff state. Building on this insight, the paper also highlights a novel channel through which information sharing becomes inefficient in beauty contest games: an agent may choose to withhold her signal to avoid the adverse consequences of her neighbors’ altered actions that would result from the increased commonality of information.
This paper is an initial step in understanding the social value of information with ex-ante heterogeneous agents. Promising avenues for future research include identifying optimal information dissemination policies, introducing endogenous information acquisition, and extending the analysis to more general linear-quadratic games.
6 Proofs
6.1 Proposition 1
@@ 6.2 Proposition 2
$\mathbb{E}[X^{2}_{i}]$ $=\mathbb{E}[((1-b_{i})\varepsilon_{i}+b_{i}\varepsilon_{y})^{2}]$
@@
$\mathbb{E}[X_{i}X_{j}]$ $=\mathbb{E}[((1-b_{i})\varepsilon_{i}+b_{i}\varepsilon_{y})((1-b_{j})\varepsilon_{j}+b_{j}\varepsilon_{y})]$
@@ 6.3 Claims in Example 1
Because $h(0,\gamma)>0$ and $h(1,\gamma)>0$ for any $\gamma\in(0,1)$, the quadratic function $h(\alpha,\gamma)$ can be strictly negative for some $\alpha\in[0,1)$ if and only if its minimum is negative. This requires two conditions to hold jointly: (i) the vertex must lie in the interval $(0,1)$, i.e., $\gamma>1/\sqrt{2}$, and (ii) the value at the vertex must be negative, i.e., $(2\gamma-1)(-4\gamma^{2}+2\gamma+1)<0$. Since $\gamma>1/\sqrt{2}$ implies that $02\gamma-1>$ 0, the two conditions are jointly equivalent to the single inequality $04\gamma^{2}-2\gamma-1>$ 0. This final inequality holds if and only if $\gamma>(1+\sqrt{5})/4$. ∎
@@
6.4 Derivation of the unique linear equilibrium under $\mathcal{I}^{\prime}$
@@
6.5 Derivation of inequality (17)
Let $U^{\prime}_{A}(\tau_{x},\tau_{y},G)$ denote Ann’s equilibrium payoff under $\mathcal{I}^{\prime}$. By the proof in Section 6.4, Ann’s and Bob’s equilibrium strategies under $\mathcal{I}^{\prime}$ are given by $a^{\prime}_{A}=(1-\gamma)(1+\gamma\alpha)y+\gamma(1-\alpha+\gamma\alpha)x_{A}$ and $a^{\prime}_{B}=(1-\gamma)y+\gamma x_{B}$, respectively. Notice that $a^{\prime}_{A}$ and $a^{\prime}_{B}$ coincide with the equilibrium strategies under $\mathcal{I}$ for agents with centralities of $1+\gamma\alpha$ and $1$, respectively. Thus, plugging $1+\gamma\alpha$ as Ann’s centrality and $1$ as Bob’s centrality into (11) yields

### v3

@@ 4.2 Aggregate Welfare
Proposition 3 implies a stark non-monotonic relationship between network connectivity and the social value of public information. To see this, I first contrast this result with the benchmark of regular networks, where all agents share the same out-degree, i.e., $d_{i}^{out}=d$ for all $i\in N$. In this environment, denser connectivity unambiguously increases the social value of public information. The left-hand side of inequality (18) simplifies to
$n(1+d-2\gamma d^{2})/(1-\gamma d)^{2},$ (20)
which is positive and strictly increasing in dd. [^11] Suppose $d_{i}^{out}=d$ for all $i\in N$. Then, the term $c_{i}^{2}(\gamma,G)-\sum_{j\neq i}g_{ij}c_{j}(\gamma,G)\left(c_{j}(\gamma,G)-2\right)$ simplifies to $(1+d-2\gamma d^{2})/(1-\gamma d)^{2}$. Differentiating this ratio with respect to $d$ yields $((1-\gamma d)+2\gamma(1-d))/(1-\gamma d)^{3}$, which is strictly positive for all $d\in[0,1)$. Thus, under the uniform interaction, the welfare effect of public information monotonically increases with network connectivity. This confirms the standard intuition that as agents become more incentivized to coordinate, the public signal becomes increasingly valuable.
However, this intuition breaks down under heterogeneous interactions, and increasing connectivity can turn a beneficial public signal into a detrimental one. The following corollary formalizes this statement.
Corollary 1.
Suppose $n\geq 3$. Then, there exist $\gamma\in(0,1)$ and two interaction networks $G$ and $G^{\prime}$ with $G\leq G^{\prime}$ such that
$\Delta W(\tau_{x},\tau_{y},G^{\prime})<0<\Delta W(\tau_{x},\tau_{y},G).$ (21)
The proof constructs the core-periphery network $G^{\prime}$ by adding weights to an empty network $G$ to concentrate connectivity on a small subset of agents, rather than densifying the network uniformly. This selective increase generates a dense core with disproportionately high centrality. By linking the other agents to this core with small but strictly positive weights, the construction exposes them to the core’s excessive volatility, while the direct benefits of coordination remain limited. Consequently, when public information is imprecise relative to private information (i.e., $\gamma$ is sufficiently high), this construction leads to a net reduction in aggregate welfare.
4.3 Endogenous Information Sharing
@@
The two-stage game is solved by backward induction. The analysis requires comparing equilibrium payoffs under public sharing with those under private retention of the signal. As derived in Section A.9, the equilibrium actions in the subgame following a choice to keep the signal private are
@@
Let $\mathcal{J}(\gamma,l,m)\equiv\{(\alpha,\beta)\in[0,1)^{2}:f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}<mf(\alpha,\beta,\gamma)+l(1-\beta(2\gamma\beta-1))\}$ be the set of interaction parameters $(\alpha,\beta)$ that admit inefficiently low equilibrium information sharing, given $\gamma$, $l$, and $m$. Because agent 1’s preference over information structures is “more aligned” with peripheral agents than with core agents, $\mathcal{J}(\gamma,l,m)$ shrinks as the number of peripheral agents relatively increases, i.e., $\mathcal{J}(\gamma,n-m^{\prime},m^{\prime})\subseteq\mathcal{J}(\gamma,n-m,m)$ for $m\leq m^{\prime}$ (see Figure 4). [^14] The formal argument is as follows. Fix $\gamma$ and take any $(\alpha,\beta)$ that satisfies $f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}$. The right-hand side of (27) is linear in $m$ with slope $f(\alpha,\beta,\gamma)-(1-\beta(2\gamma\beta-1))$ by substituting l=nml=n-m. This slope is bounded from above by $f(\alpha,\beta,\gamma)-1+2\gamma\beta^{2}-\beta$ $\leq f(\alpha,\beta,\gamma)-1+2\gamma\beta-\beta$ $\leq f(\alpha,\beta,\gamma)-1+2\gamma\beta-\gamma\beta$ $=f(\alpha,\beta,\gamma)-1+\gamma\beta$ $<0.$ The last strict inequality follows from the initial assumption. Thus, the right-hand side of (27) strictly decreases with $m$ whenever $f(\alpha,\beta,\gamma)<(1-\beta\gamma)^{2}$, which proves the claim.
@@
5 Discussion
5.1 Marginal Value of Public Information
While the preceding analysis identifies conditions for discrete welfare losses, a more calibrated policy requires characterizing the marginal value of transparency. This section examines the sensitivity of aggregate welfare to a local increase in the public signal precision $\tau_{y}$.
Recall that, based on the equilibrium characterization, aggregate welfare is simplified to
$W(\tau_{x},\tau_{y},G)=\tau_{x}^{-1}(1-\gamma)S(\gamma)+C,$ (28)
where $C$ is a constant independent of $\tau_{y}$, and $S(\gamma)$ captures the network-dependent welfare component:
$S(\gamma)\equiv\sum_{i\in N}c_{i}(\gamma,G)\left((1-d_{i}^{in})c_{i}(\gamma,G)+2d_{i}^{in}\right).$ (29)
Increasing transparency creates a tension between signal reliance and network feedback. On one hand, a higher $\tau_{y}$ increases the baseline weight $1-\gamma$ placed on the public signal. This scaling effect amplifies the distortion captured by $S(\gamma)$. On the other hand, it reduces the discount factor $\gamma$ governing centrality. This dampening effect shrinks the centrality vector, attenuating the magnitude of the distortion. The centrality sensitivity vector $\mathbf{c}^{\mathbf{c}}(\gamma,G)\equiv(I-\gamma G)^{-1}\mathbf{c}(\gamma,G)$ captures the sensitivity of centrality to variations in $\gamma$, as the derivative of centrality vector with respect to $\gamma$ is proportional to the difference $\mathbf{c}^{\mathbf{c}}(\gamma,G)-\mathbf{c}(\gamma,G)$. [^15] Differentiating $\mathbf{c}=(I-\gamma G)^{-1}\mathbf{1}$ yields $\partial\mathbf{c}/\partial\gamma=(I-\gamma G)^{-1}G\mathbf{c}$. Using the identity $\mathbf{c}=\mathbf{1}+\gamma G\mathbf{c}$, I substitute $G\mathbf{c}=\gamma^{-1}(\mathbf{c}-\mathbf{1})$ to obtain $\partial\mathbf{c}/\partial\gamma=\gamma^{-1}\left((I-\gamma G)^{-1}\mathbf{c}-(I-\gamma G)^{-1}\mathbf{1}\right)$. By definition, this simplifies to $\gamma^{-1}(\mathbf{c}^{\mathbf{c}}-\mathbf{c})$.
Proposition 5.
The marginal value of public information is negative, i.e., $\partial W/\partial\tau_{y}<0$, if and only if
$S(\gamma)<(1-\gamma)S^{\prime}(\gamma),$ (30)
where
$S^{\prime}(\gamma)=2\gamma^{-1}\sum_{i\in N}\left(c^{\mathbf{c}}_{i}(\gamma,G)-c_{i}(\gamma,G)\right)\left((1-d_{i}^{in})c_{i}(\gamma,G)+d_{i}^{in}\right).$ (31)
Inequality (30) imposes a more stringent condition than the one for a discrete welfare loss identified in Proposition 3. While a discrete loss arises whenever the aggregate distortion is negative, a marginal welfare reduction requires this distortion to be severe enough to overwhelm the dampening effect. [^16] Note that when the network exhibits a discrete welfare loss (i.e., $S(\gamma)<0$), the sensitivity term $S^{\prime}(\gamma)$ is typically negative. This occurs because agents who have high centrality and high in-degree are also the most sensitive to changes in $\gamma$, implying a large gap between their centrality and its sensitivity (i.e., $c_{i}^{\mathbf{c}}(\gamma,G)-c_{i}(\gamma,G)$). It is therefore possible for a network to exhibit a discrete welfare loss yet benefit from marginal increases in transparency if the dampening effect is sufficiently strong. This distinction vanishes in the limit where public information becomes negligible (i.e., $\gamma\to 1$). In this case, the sensitivity term on the right-hand side of (30) converges to zero, reducing the condition to $S(1)<0$. This confirms that an interaction network generating a discrete welfare loss must also make the initial introduction of public information marginally detrimental.
5.2 Alternative Payoff Specification
This section generalizes the payoff function to disentangle the intrinsic motive for coordination from the structural constraints of the interaction network. Agent $i$’s payoff function is given by
$u_{i}(a,\theta)=-\left(1-r_{i}\right)(a_{i}-\theta)^{2}-r_{i}\sum_{j\neq i}g_{ij}(a_{i}-a_{j})^{2},$ (32)
The parameter $r_{i}\in(0,1)$ captures the relative intensity of coordination motives against adaptation motives. The corresponding first-order condition yields
$a_{i}=\left(1-\tilde{d}_{i}^{out}\right)\mathbb{E}_{i}[\theta]+\sum_{j\neq i}\tilde{g}_{ij}\mathbb{E}_{i}[a_{j}],$ (33)
where $\tilde{g}_{ij}$ denotes the $(i,j)$-th element of the normalized interaction network $\tilde{G}$, defined by
$\tilde{g}_{ij}\equiv\frac{r_{i}g_{ij}}{1-r_{i}+r_{i}d_{i}^{out}},$ (34)
and $\tilde{d}_{i}^{out}\equiv\sum_{j\neq i}\tilde{g}_{ij}$ is the normalized out-degree of agent $i$.
The normalized interaction weight $\tilde{g}_{ij}$ is strictly increasing in both the pairwise link strength $g_{ij}$ and the coordination intensity $r_{i}$, but strictly decreasing in the out-degree $d_{i}^{out}$. [^17] This captures a “dilution effect”: as an agent’s neighborhood expands, the relative influence of any single peer $j$ diminishes, effectively spreading the agent’s coordination capacity across a wider set of interactions. Because equation (33) takes the same form as the best-response condition (2) in the baseline model with $\tilde{G}$ replacing $G$, and $\tilde{G}$ is row-substochastic, the game admits a unique linear equilibrium characterized by the Katz-Bonacich centrality on $\tilde{G}$. This equivalence allows the welfare analysis of the baseline model to be naturally extended to this generalized setting.
Proposition 6.
The welfare effect of public information provision is negative, i.e., $\Delta W(\tau_{x},\tau_{y},G)<0$, if and only if
$\sum_{i\in N}c_{i}(\gamma,\tilde{G})\left(\left(K_{i}-d_{i}^{in,r}\right)c_{i}(\gamma,\tilde{G})+2d_{i}^{in,r}\right)<0,$ (35)
where $K_{i}\equiv 1-r_{i}+r_{i}d_{i}^{out}$ and $d_{i}^{in,r}\equiv\sum_{j\neq i}r_{j}g_{ji}.$
In inequality (35), the term $K_{i}$ represents agent $i$’s total payoff weight, capturing the combined magnitude of losses from misaligned adaptation and coordination. The term $d_{i}^{in,r}$ measures agent $i$’s effective influence, derived by weighting the coordination motives directed at $i$ by the intensities of her neighbors. Thus, aggregate welfare decreases when disproportionately central agents exert high effective influence relative to their total payoff weight.
This result implies that neither heterogeneity in directionality (i.e., $r_{i}=r$ and $d_{i}^{out}=1$, but $g_{ij}$ varies) nor heterogeneity in intensity (i.e., $g_{ij}$ is uniform, but $r_{i}$ varies) is sufficient to generate negative welfare effects in isolation. Instead, the detrimental effect requires the interplay of these two factors. The following corollary formalizes this necessity.
Corollary 2.
The welfare effect of public information provision is positive if either (1) $r_{i}=r$ and $d_{i}^{out}=1$ for all $i\in N$, or (2) $g_{ij}=1/(n-1)$ for all $i\neq j$.
5.3 Efficient Use of Information
To analyze how the welfare losses identified above relate to agents’ use of information, I characterize the efficient use of information following the approach of Angeletos and Pavan (2007). This efficient benchmark is defined as the strategy profile that maximizes aggregate welfare, subject to the constraint that agents must condition their actions solely on their own decentralized information sets. By keeping the informational friction constant while internalizing payoff interdependencies, this comparison isolates the specific inefficiency driven by the gap between private and social incentives.
In my framework, the efficient strategies are equivalent to the equilibria of a beauty contest game played on a modified network. Specifically, the following first-order condition characterizes the efficient strategies:
$a_{i}=\left(1-d_{i}^{F}\right)\mathbb{E}_{i}[\theta]+\sum_{j\neq i}g_{ij}^{F}\mathbb{E}_{i}[a_{j}],$ (36)
where $d_{i}^{F}\equiv\sum_{j\neq i}g_{ij}^{F}$ and $G^{F}=[g_{ij}^{F}]$ is the fictitious interaction network defined by
$g_{ij}^{F}\equiv\frac{g_{ij}+g_{ji}}{1+d_{i}^{in}}.$ (37)
This fictitious network captures the wedge between private and social coordination motives. The numerator strengthens bilateral coordination incentives by internalizing both sides’ payoffs, while the denominator scales these bilateral motives down more strongly for high in-degree agents than for low in-degree ones. Using this equivalence, Proposition 7 presents a unique linear efficient strategy.
Proposition 7.
There exists a unique linear efficient strategy profile in which agent $i$’s strategy is
$a_{i}^{F}=(1-\gamma)c_{i}(\gamma,G^{F})y+(1-(1-\gamma)c_{i}(\gamma,G^{F}))x_{i},$ (38)
where $c_{i}(\gamma,G^{F})$ is $i$’s Katz-Bonacich centrality in $G^{F}$.
This result contrasts with the findings of Angeletos and Pavan (2007). In their analysis of symmetric settings, they show that when the social degree of coordination exceeds the private one, agents under-respond to public information in equilibrium. [^18] In Angeletos and Pavan (2007), the private degree of coordination represents the weight assigned to the average action in the equilibrium first-order condition. In contrast, the social degree of coordination represents the corresponding weight in the planner’s first-order condition. In contrast, in the present framework, the direction of inefficiency is governed by network topology; in particular, agents over-respond to public information in equilibrium whenever their private centrality exceeds their social centrality, $c_{i}(\gamma,G)>c_{i}(\gamma,G^{F})$.
To illustrate the locus of inefficiency, consider Example 1 of Ann, Bob, and Carol. Figure 5 presents the corresponding fictitious network. In this fictitious network, Bob’s weight on Ann increases with $\alpha$, reflecting the planner’s internalization of the unilateral externality Bob exerts on Ann. Conversely, Bob’s weight on Carol decreases with $\alpha$, thereby weakening the coordination loop between Bob and Carol as Ann’s dependence on Bob intensifies.
[…]
Figure 5: The fictitious interaction network corresponding to Figure 1.
This structural transformation identifies Bob as the primary source of inefficiency. Consider a numerical specification with a strong core ($\beta=0.95$), a weakly dependent periphery ($\alpha=0.2$), and high relative precision of private signals ($\gamma=0.95$). Under these parameters, Bob’s centrality strictly decreases from a private level of $c_{B}(\gamma,G)\approx 10.26$ to a social level of $c_{B}(\gamma,G^{F})\approx 9.35$. [^19] In the fictitious network, the weights incident to Bob are $g_{BA}^{F}=\alpha/(1+\alpha+\beta)$ and $g_{BC}^{F}=2\beta/(1+\alpha+\beta)$, while the return links are $g_{AB}^{F}=\alpha$ from Ann and $g_{CB}^{F}=2\beta/(1+\beta)$ from Carol. Bob’s centrality satisfies the recursive condition $c_{B}=1+\gamma g_{BA}^{F}c_{A}+\gamma g_{BC}^{F}c_{C}$. Substituting the expressions for neighbors’ centralities, $c_{A}=1+\gamma\alpha c_{B}$ and $c_{C}=1+\gamma g_{CB}^{F}c_{B}$, into Bob’s equation yields $c_{B}=1+\gamma g_{BA}^{F}(1+\gamma\alpha c_{B})+\gamma g_{BC}^{F}(1+\gamma g_{CB}^{F}c_{B}).$ Grouping the terms with $c_{B}$ on the left-hand side and solving yields the closed-form solution: $c_{B}(\gamma,G^{F})=\frac{1+\gamma(g_{BA}^{F}+g_{BC}^{F})}{1-\gamma^{2}(\alpha g_{BA}^{F}+g_{BC}^{F}g_{CB}^{F})}.$ For parameters $\beta=0.95,\alpha=0.2$, and $\gamma=0.95$, $c_{B}(\gamma,G^{F})=808275/86408$, which is strictly lower than the private centrality $c_{B}(\gamma,G)=400/39$. Thus, Bob over-responds to public information relative to the efficient benchmark. This localized excess coordination creates a negative externality for Ann, mirroring the classic tension between private and social incentives highlighted by Morris and Shin (2002).
@@ 6 Conclusion
This paper investigates how the structure of interaction networks determines the welfare effects of public information. I show that the sign of the welfare effect depends on the interplay between Katz-Bonacich centrality and aggregate influence (i.e., in-degree). In the tiered core-periphery architectures typical of OTC markets, a small, dense core is sufficiently central and influential that its excessive responsiveness to public signals generates a volatility externality, outweighing the coordination benefits for the periphery. This mechanism also uncovers a strategic friction in information sharing that leads to the under-provision of public information: agents may rationally withhold signals to avoid exposure to the volatility induced by their highly connected peers.
These findings imply that the value of transparency is non-monotonic in network connectivity. While uniform connectivity enhances the value of public information, the emergence of dominant hubs can reverse it. Future research could investigate how these forces shape endogenous network formation, particularly whether agents strategically limit connections to mitigate the volatility externality identified here.
Appendix: Proofs
A.1 Proposition 1
@@ A.2 Proposition 2
$\mathbb{E}[X^{2}_{i}]$ $=\mathbb{E}[((1-b_{i})\varepsilon_{i}+b_{i}\varepsilon_{0})^{2}]$
@@
$\mathbb{E}[X_{i}X_{j}]$ $=\mathbb{E}[((1-b_{i})\varepsilon_{i}+b_{i}\varepsilon_{0})((1-b_{j})\varepsilon_{j}+b_{j}\varepsilon_{0})]$
@@ A.3 Claims in Example 1
Because $h(0,\gamma)>0$ and $h(1,\gamma)>0$ for any $\gamma\in(0,1)$, the quadratic function $h(\alpha,\gamma)$ can be strictly negative for some $\alpha\in[0,1)$ if and only if its minimum is negative. This requires two conditions to hold jointly: (i) the vertex must lie in the interval $(0,1)$, i.e., $\gamma>1/\sqrt{2}$, and (ii) the value at the vertex must be negative, i.e., $(2\gamma-1)(-4\gamma^{2}+2\gamma+1)<0$. Since $\gamma>1/\sqrt{2}$ implies that $2\gamma-1>0$, the two conditions are jointly equivalent to the single inequality $4\gamma^{2}-2\gamma-1>0$. This final inequality holds if and only if $\gamma>(1+\sqrt{5})/4$. ∎
@@
A.4 Derivation of the unique linear equilibrium under $\mathcal{I}^{\prime}$
@@
A.5 Derivation of inequality (17)
Let $U^{\prime}_{A}(\tau_{x},\tau_{y},G)$ denote Ann’s equilibrium payoff under $\mathcal{I}^{\prime}$. By the proof in Section A.4, Ann’s and Bob’s equilibrium strategies under $\mathcal{I}^{\prime}$ are given by $a^{\prime}_{A}=(1-\gamma)(1+\gamma\alpha)y+\gamma(1-\alpha+\gamma\alpha)x_{A}$ and $a^{\prime}_{B}=(1-\gamma)y+\gamma x_{B}$, respectively. Notice that $a^{\prime}_{A}$ and $a^{\prime}_{B}$ coincide with the equilibrium strategies of agents with centralities of $1+\gamma\alpha$ and $1$ under $\mathcal{I}$, respectively. Thus, plugging $1+\gamma\alpha$ as Ann’s centrality and $1$ as Bob’s centrality into (11) yields